# MGPATH: Vision-Language Model with Multi-Granular Prompt Learning for Few-Shot WSI Classification

### v3

@@ A.1 Other hyper-parameters
For all experiments, we trained MGPath with the Adam optimizer with a learning rate of $9\times 10^{-6}$ and a weight decay of $1\times 10^{-5}$ to fine-tune all versions of MGPath presented in Tables  1 and  2. The training process was conducted for a maximum of 200 epochs, with a batch size set to 1. The best checkpoints are picked based validation performance with F1 score.
@@ A.2 Baseline Setups
TCGA-BRCA: The baselines in Table  1 are sourced from the MSCPT Han et al. (2024) paper, where various methods are evaluated using two backbones: Vision Transformer (ViT)  Alexey (2020) from the CLIP model (top section of Table  1) and PLIP  Huang et al. (2023) (bottom section of Table  1). In this context, we introduce three variations of MGPath - ViT, PLIP, and GigaPath-PLIP (abbreviated as PLIP-G), where all versions utilize frozen vision and text encoders.
TCGA-RCC &\&& TCGA-NSCLC: The baselines in Table  2 are adapted from ViLa-MIL Shi et al. (2024) where methods employ ResNet-50 from the CLIP model as the primary backbone. We present MGPath results using three architectures: ResNet-50, PLIP, and PLIP-G. With ResNet-50, we follow the ViLa-MIL approach by training the text encoder and reporting performance for this setup. To assess efficiency, we provide the total parameter counts for both ViLa-MIL and MGPath, considering scenarios with frozen backbones and trainable text encoders. For PLIP and PLIP-G, all visual and text encoders are kept frozen.
CONCH &\&& QUILT: We download the pre-trained weights of these foundation models and adapt them for zero-shot evaluation on TCGA datasets following the authors’ guidelines from  Lu et al. (2024), which randomly sample 75 samples for each class. For few-shot settings, since official implementations are not provided, we initialize the models with their pre-trained weights and allow fully fine-tune the text encoder and evaluate on the same subsets that we use for other baselines. While CONCH provides prompts for the datasets in its publication, QUILT does not. Therefore, we fine-tune the model using CONCH’s prompts and our own generated prompts for QUILT.
@@
MGPath (GAT CONV) 77.2$\pm$1.3 70.9$\pm$2.0 71.0$\pm$2.1
MGPath (GIN) 77.1$\pm$2.9 69.8$\pm$3.9 69.9$\pm$4.0
MGPath (GCN) 75.1$\pm$2.9 67.6$\pm$2.5 67.1$\pm$2.8
Table 8: AUC performance comparison over epochs for PLIP (blue) and PLIP combined with GigaPath (red). GigaPath significantly enhances the AUC, achieving more stable and higher values, particularly in the early epochs.
Appendix B Impact of PLIP enhanced Prov-GigaPath
@@ Appendix C Ablation Study on Message Passing Networks
In Table 8, we evaluate the performance of MGPath (CLIP) using the Graph Attention Network (GAT-CONV) against alternatives like the Graph Isomorphism Network (GIN) Xu et al. (2019b) and the Graph Convolutional Network (GCN) Kipf and Welling (2017). The results show that MGPath (GIN) achieves comparable performance to MGPath (GAT-CONV), however, with higher variance. In contrast, MGPath (GAT-CONV) significantly outperforms the GCN-based version, likely due to GAT’s ability to dynamically assign attention weights to neighboring image patches, enabling it to prioritize the most relevant neighbors for each node.
Appendix D Additional Details on Optimal Transport Distance
The following paragraphs will provide detailed information on the implementation of (un-balanced) optimal transport (OT) Villani et al. (2009); Peyré et al. (2019) and specifically the alignment of prompt-guided visual-text distances in MGPath.
D.1 OT Formulation and Efficient Solver
Given two set of feature embeddings $\bm{F}=\left\{\bm{f}_{i}|_{i=1}^{M}\right\}\in\mathbb{R}^{M\times d}$ and $\bm{G}=\left\{\bm{g}_{j}|_{j=1}^{N}\right\}\in\mathbb{R}^{N\times d}$, we can represent them as two discrete distributions $\bm{\mu}$ and $\bm{\nu}$ by:
$\bm{\mu}=\sum^{M}_{i=1}p_{i}\delta_{\bm{f}_{i}},\quad\bm{\nu}=\sum^{N}_{j=1}q_{j}\delta_{\bm{g}_{j}},$ (12)
where $\delta_{\bm{f}_{i}}$ and $\delta_{\bm{g}_{j}}$ represent Dirac delta functions centered at $\bm{F}$ and $\bm{G}$, respectively and the weights are elements of the marginal $\bm{p}=\{p_{i}\}^{M}_{i=1}$ and $\bm{q}=\{q_{i}\}^{N}_{j=1}$ and can be selected as the uniform weight with $\sum_{i=1}^{M}p_{i}=1$, $\sum_{j=1}^{N}q_{j}=1$.
Then we can compute the distance between $\bm{F}$ and $\bm{G}$ through $\bm{\mu}$ and $\bm{\nu}$ (Eq.(9)) as
$d_{\mathrm{OT}}(\bm{\mu},\bm{\nu})=\langle\bm{T}^{\ast},\bm{C}\rangle.$ (13)
where
$\bm{T}^{\ast}=\underset{\bm{T}\in\mathbb{R}^{MXN}}{\arg{\min}}\sum^{M}_{i=1}\sum^{N}_{j=1}\bm{T}_{ij}\bm{C}_{ij}$
@@ D.1 OT Formulation and Efficient Solver
with $\bm{C}\in\mathbb{R}^{M\times N}$ is the cost matrix which measures the distance between $\bm{f}_{i}\in\bm{\mu}$ and $\bm{g}_{j}\in\bm{\nu}$.
Because directly solving Eq (14) is high-computational costs ($O(n^{3}\log n)$ with $n$ proportional to $M$ and $N$), Sinkhorn algorithm Cuturi (2013) is proposed to approximate solution by solving a regularized problem:
$\bm{T}^{\ast}=\underset{\bm{T}\in\mathbb{R}^{MXN}}{\arg{\min}}\sum^{M}_{i=1}\sum^{N}_{j=1}\bm{T}_{ij}\bm{C}_{ij}-\lambda H(\bm{T})$
[…]
where $H(\bm{T})$ = $\sum_{ij}\bm{T}_{ij}\log\bm{T}_{ij}$ be an entropy function and $\lambda>0$ is the regularization parameter. The optimization problem in Eq.  (15) is strictly convex, allowing us to achieve a solution efficiently with fewer iterations as outlined below:
$\bm{T}^{\ast}=\text{diag}(\bm{a}^{t})\exp(-\mathbb{C}/\lambda)\text{diag}(\bm{b}^{t})$ (16)
where $t$ is the iteration and $\bm{a}^{t}=\bm{\mu}/\exp(-\mathbb{C}/\lambda)\bm{b}^{t-1}$ and $\bm{b}^{t}=\bm{\nu}/\exp(-\mathbb{C}/\lambda)\bm{a}^{t}$, with the initialization on $\bm{b}^{0}=\bm{1}$. In our experiments, we used $t=100$ and $\lambda=0.1$ based on validation performance.
D.2 Relaxed Marginal Constraints with Unbalanced Optimal Transport
Due to strict marginal constraints in Eq (14), optimal transport may be unrealistic in real-world scenarios where data distributions are noisy, incomplete, or unbalanced. The Unbalanced Optimal Transport (UoT) Chizat et al. (2018); Liero et al. (2018) addresses this challenge by relaxing the marginal constraints, allowing for partial matching through penalties on mass creation or destruction. In particular, UoT solves
$\bm{T}^{\ast}=\underset{\bm{T}\in\mathbb{R}^{MXN}}{\arg{\min}}\sum^{M}_{i=1}\sum^{N}_{j=1}\bm{T}_{ij}\bm{C}_{ij}-\lambda H(\bm{T})$ (17)
$+\ \rho_{1}\mathrm{KL}(\bm{T}\bm{1}^{N}||\bm{\mu})+\ \rho_{1}\mathrm{KL}(\bm{T}^{\top}\bm{1}^{M}||\bm{\nu})$
here, $\rho_{1}$ and $\rho_{2}$ represent the marginal regularization parameters, and $\mathrm{KL}(\bm{P}||\bm{Q})$ denotes the Kullback-Leibler divergence between two positive vectors. Similar to the classical OT formulation, there are solvers based on the Sinkhorn algorithm that can address Eq. (D.2) Pham et al. (2020). However, these solvers typically require more iteration steps to converge to optimal solutions due to the added complexity introduced by the relaxed marginal constraints.
Appendix E Unbalance Optimal Transport (UoT)
Table 9: MGPath performance and running time (in second) comparison between OT and UoT.
@@ Appendix E Unbalance Optimal Transport (UoT)
MGPath (OT, 4 text prompts) 76.2$\pm$2.2 69.0$\pm$3.5 69.3$\pm$2.8 1482
MGPath (UoT, 4 text prompts) 77.0$\pm$1.8 70.2±$\pm 3.4$ 70.4±$\pm 3.3$ 3260
@@
MGPath (OT, 4 text prompts) 92.1$\pm$2.8 76.5$\pm$5.2 81.7$\pm$2.9 1451
MGPath (UoT, 4 text prompts) 92.8$\pm$2.4 76.8$\pm$4.7 82.4 $\pm$ 2.4 3049
To conduct a comparative evaluation of the performance of MGPath using optimal transport versus unbalanced optimal transport (Section D.2) given the more flexible constraints in UoT, we conducted an additional experiment. To be specific, we test on the TCGA-NSCLC and TCGA-RCC datasets with the CLIP architecture (ResNet-50) using a 4-text-prompt setting. Table  9 presents our findings where the running time is computed as seconds of average across five-folds. The results show that UoT outperforms OT with an approximate 1% improvement across all metrics. However, UoT is approximately 2 times slower than OT. This increase is attributed to the added flexibility and complexity introduced by relaxing the marginal constraints in the UoT formulation. Given this trade-off, we choose OT as the main distance in MGPath and leave the UoT version for further evaluation. It is also important to know that our OT formulation leverages approximate solutions through the regularized formulation (Eq.,(15)) and produces smoothed optimal mappings $\bm{T}^{\ast}$, which can implicitly help the model adapt to perturbations like UoT.

### v5

@@ A.1 Other hyper-parameters
For all experiments, we trained MGPATH with the Adam optimizer with a learning rate of $9\times 10^{-6}$ and a weight decay of $1\times 10^{-5}$ to fine-tune all versions of MGPATH presented in Tables  4 and  1. The training process was conducted for a maximum of 200 epochs, with a batch size set to 1. The best checkpoints are picked based validation performance with F1 score.
@@ A.2 Baseline Setups
TCGA-BRCA: The baselines in Table  4 are sourced from the MSCPT (Han et al., 2024) paper, where various methods are evaluated using two backbones: Vision Transformer (ViT)  (Alexey, 2020) from the CLIP model (top section of Table  4) and PLIP  (Huang et al., 2023) (bottom section of Table  4). In this context, we introduce three variations of MGPATH(ViT), MGPATH(PLIP), and MGPATH(PLIP-G), where all versions utilize frozen vision and text encoders.
TCGA-RCC &\& TCGA-NSCLC: The baselines in Table  1 are adapted from ViLa-MIL (Shi et al., 2024) where methods employ ResNet-50 from the CLIP model as the primary backbone. We present the results using three architectures: MGPATH(CLIP), MGPATH(PLIP), and MGPATH(PLIP-G). With ResNet-50, we follow the ViLa-MIL approach by training the text encoder and reporting performance for this setup. To assess efficiency, we provide the total parameter counts for both ViLa-MIL and MGPATH, considering scenarios with frozen backbones and trainable text encoders. For MGPATH(PLIP) and MGPATH(PLIP-G), all visual and text encoders are kept frozen.
CONCH &\& QUILT: We download the pre-trained weights of these foundation models and adapt them for zero-shot evaluation on TCGA datasets following the authors’ guidelines from  (Lu et al., 2024), which randomly sample 75 samples for each class. For few-shot settings, since official implementations are not provided, we initialize the models with their pre-trained weights and allow fully fine-tune the text encoder and evaluate on the same subsets that we use for other baselines. While CONCH provides prompts for the datasets in its publication, QUILT does not. Therefore, we fine-tune the model using CONCH’s prompts and our own generated prompts for QUILT.
Appendix B Impact of PLIP enhanced Prov-GigaPath
@@
[…]
Table 9: AUC performance comparison over epochs for PLIP (blue) MGPATH(PLIP-G) (red). MGPATH(PLIP-G) achieved achieving more stable and higher values, particularly in the early epochs.
@@ Appendix C Ablation Study on Message Passing Networks
In Table 9, we evaluate the performance of MGPATH(CLIP) using the Graph Attention Network (GAT-CONV)against alternatives like the Graph Isomorphism Network (GIN) (Xu et al., 2019b) and the Graph Convolutional Network (GCN) (Kipf & Welling, 2017). The results show that MGPATH (GIN) achieves comparable performance to MGPATH(CLIP)(GAT-CONV), however, with higher variance. In contrast, MGPATH(CLIP)(GAT-CONV) significantly outperforms the GCN-based version, likely due to GAT’s ability to dynamically assign attention weights to neighboring image patches, enabling it to prioritize the most relevant neighbors for each node.
Appendix D Ablation Studies on Data Augmentation
To evaluate the effect of data augmentation on model performance, we conducted experiments comparing MGPATH(PLIP-G) with and without augmentation. In addition, we trained a variant, MGPATH(PLIP-G)(COS), where the optimal transport module in the MGPATH(PLIP-G) architecture was replaced by cosine similarity. All experiments were carried out on the TCGA-NSCLC dataset under the 16-shot setting.
As illustrated in Table 10, the cosine similarity configuration exhibits unstable behavior, with a sharp decline in F1 and ACC when data augmentation is applied, indicating difficulty in adapting to heterogeneous data in the training data. In contrast, the optimal transport configuration not only achieves the best overall results but also remains consistently strong across augmented and non-augmented settings.
Table 10: Impact of data augmentation on the performance of MGPATH using cosine similarity and optimal transport configurations on the TCGA-NSCLC dataset.
Method TCGA-NSCL
AUC F1 ACC
MGPATH(PLIP-G)(COS) 91.13 $\pm$ 3.20 82.88 $\pm$ 3.55 83.05 $\pm$ 3.53
w/o augmentation 92.92 $\pm$ 3.88 74.35 $\pm$ 23.40 78.17 $\pm$ 15.93
MGPATH(PLIP-G) 93.95 $\pm$ 2.71 88.53 $\pm$ 4.29 88.63 $\pm$ 4.26
w/o augmentation 91.94 $\pm$ 5.33 84.49 $\pm$ 6.02 84.67 $\pm$ 5.98
Appendix E Training and Inference Efficiency
To assess the computational efficiency of MGPATH(PLIP-G) based on OT and cosine similarity, MGPATH(PLIP-G)(OT) and MGPATH(PLIP-G)(COS), respectively. We conducted an additional experiments. To concreat, we evaluate training and inference using the metrics Throughput (samples/sec) on the dataset TCGA-NSCLC with text prompts generated by Mistral Medium 3  (Mistral AI, 2024) in 16-shots setting.
Table 11 shows that although the optimal transport configuration requires more computation and yields lower training and inference throughput compared to cosine similarity, it consistently achieves higher accuracy. This highlights the strength of optimal transport in improving model generalization, as it is able to leverage complex data distributions more effectively despite the additional computational overhead.
Table 11: Comparison of average Training and Inference Throughput (samples/sec) on a single NVIDIA A100 GPU with only batch size 1 for MGPATH(PLIP-G)(OT) and MGPATH(PLIP-G)(COS).
Configuration Training Inference
@@ Appendix E Training and Inference Efficiency
MGPATH(PLIP-G)(COS) 16.00 0.29 31.59 0.32 83.05 3.53
MGPATH(PLIP-G) 10.67 2.52 19.05 5.59 88.63 4.26
Appendix F Unbalance Optimal Transport (UoT)
To conduct a comparative evaluation of the performance of MGPATH using optimal transport versus unbalanced optimal transport (Section G.2) given the more flexible constraints in UoT, we conducted an additional experiment. To be specific, we test on the TCGA-NSCLC and TCGA-RCC datasets with MGPATH(CLIP) using a 4-text-prompt setting. Table  12 presents our findings where the running time is computed as seconds of average across five-folds. The results show that UoT outperforms OT with an approximate 1% improvement across all metrics. However, UoT is approximately 2 times slower than OT. This increase is attributed to the added flexibility and complexity introduced by relaxing the marginal constraints in the UoT formulation. Given this trade-off, we choose OT as the main distance in MGPATH and leave the UoT version for further evaluation. It is also important to know that our OT formulation leverages approximate solutions through the regularized formulation (Eq.,(15)) and produces smoothed optimal mappings $\boldsymbol{T}^{\ast}$, which can implicitly help the model adapt to perturbations like UoT.
Table 12: MGPATH(CLIP) performance and running time (in second) comparison between OT and UoT.
@@ Appendix F Unbalance Optimal Transport (UoT)
[…]
@@
[…]
Appendix G Additional Details on Optimal Transport Distance
The following paragraphs will provide detailed information on the implementation of (un-balanced) optimal transport (OT)  (Villani et al., 2009; Peyré et al., 2019) and specifically the alignment of prompt-guided visual-text distances in MGPATH.
G.1 OT Formulation and Efficient Solver
Given two set of feature embeddings $\boldsymbol{F}=\left\{\boldsymbol{f}_{i}|_{i=1}^{M}\right\}\in\mathbb{R}^{M\times d}$ and $\boldsymbol{G}=\left\{\boldsymbol{g}_{j}|_{j=1}^{N}\right\}\in\mathbb{R}^{N\times d}$, we can represent them as two discrete distributions $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$ by:
$\boldsymbol{\mu}=\sum^{M}_{i=1}p_{i}\delta_{\boldsymbol{f}_{i}},\quad\boldsymbol{\nu}=\sum^{N}_{j=1}q_{j}\delta_{\boldsymbol{g}_{j}},$ (12)
where $\delta_{\boldsymbol{f}_{i}}$ and $\delta_{\boldsymbol{g}_{j}}$ represent Dirac delta functions centered at $\boldsymbol{F}$ and $\boldsymbol{G}$, respectively and the weights are elements of the marginal $\boldsymbol{p}=\{p_{i}\}^{M}_{i=1}$ and $\boldsymbol{q}=\{q_{i}\}^{N}_{j=1}$ and can be selected as the uniform weight with $\sum_{i=1}^{M}p_{i}=1$, $\sum_{j=1}^{N}q_{j}=1$.
Then we can compute the distance between $\boldsymbol{F}$ and $\boldsymbol{G}$ through $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$ (Eq.(9)) as
$d_{\mathrm{OT}}(\boldsymbol{\mu},\boldsymbol{\nu})=\langle\boldsymbol{T}^{\ast},\boldsymbol{C}\rangle.$ (13)
where
$\boldsymbol{T}^{\ast}=\underset{\boldsymbol{T}\in\mathbb{R}^{MXN}}{\arg{\min}}\sum^{M}_{i=1}\sum^{N}_{j=1}\boldsymbol{T}_{ij}\boldsymbol{C}_{ij}$
$\textrm{s.t.}\quad\boldsymbol{T}\boldsymbol{1}^{N}=\boldsymbol{\mu},\quad\boldsymbol{T}^{\top}\boldsymbol{1}^{M}=\boldsymbol{\nu}.$ (14)
with $\boldsymbol{C}\in\mathbb{R}^{M\times N}$ is the cost matrix which measures the distance between $\boldsymbol{f}_{i}\in\boldsymbol{\mu}$ and $\boldsymbol{g}_{j}\in\boldsymbol{\nu}$.
Because directly solving Eq (14) is high-computational costs ($O(n^{3}\log n)$ with $n$ proportional to $M$ and $N$), Sinkhorn algorithm (Cuturi, 2013) is proposed to approximate solution by solving a regularized problem:
$\boldsymbol{T}^{\ast}=\underset{\boldsymbol{T}\in\mathbb{R}^{MXN}}{\arg{\min}}\sum^{M}_{i=1}\sum^{N}_{j=1}\boldsymbol{T}_{ij}\boldsymbol{C}_{ij}-\lambda H(\boldsymbol{T})$
$\textrm{s.t.}\quad\boldsymbol{T}\boldsymbol{1}^{N}=\boldsymbol{\mu},\quad\boldsymbol{T}^{\top}\boldsymbol{1}^{M}=\boldsymbol{\nu}.$ (15)
where $H(\boldsymbol{T})$ = $\sum_{ij}\boldsymbol{T}_{ij}\log\boldsymbol{T}_{ij}$ be an entropy function and $\lambda>0$ is the regularization parameter. The optimization problem in Eq.  (15) is strictly convex, allowing us to achieve a solution efficiently with fewer iterations as outlined below:
$\boldsymbol{T}^{\ast}=\text{diag}(\boldsymbol{a}^{t})\exp(-\mathbb{C}/\lambda)\text{diag}(\boldsymbol{b}^{t})$ (16)
where $t$ is the iteration and $\boldsymbol{a}^{t}=\boldsymbol{\mu}/\exp(-\mathbb{C}/\lambda)\boldsymbol{b}^{t-1}$ and $\boldsymbol{b}^{t}=\boldsymbol{\nu}/\exp(-\mathbb{C}/\lambda)\boldsymbol{a}^{t}$, with the initialization on $\boldsymbol{b}^{0}=\boldsymbol{1}$. In our experiments, we used $t=100$ and $\lambda=0.1$ based on validation performance.
G.2 Relaxed Marginal Constraints with Unbalanced Optimal Transport
Due to strict marginal constraints in Eq (14), optimal transport may be unrealistic in real-world scenarios where data distributions are noisy, incomplete, or unbalanced. The Unbalanced Optimal Transport (UoT)  (Chizat et al., 2018; Liero et al., 2018) addresses this challenge by relaxing the marginal constraints, allowing for partial matching through penalties on mass creation or destruction. In particular, UoT solves
$\boldsymbol{T}^{\ast}=\underset{\boldsymbol{T}\in\mathbb{R}^{MXN}}{\arg{\min}}\sum^{M}_{i=1}\sum^{N}_{j=1}\boldsymbol{T}_{ij}\boldsymbol{C}_{ij}-\lambda H(\boldsymbol{T})$ (17)
$+\ \rho_{1}\mathrm{KL}(\boldsymbol{T}\boldsymbol{1}^{N}||\boldsymbol{\mu})+\ \rho_{1}\mathrm{KL}(\boldsymbol{T}^{\top}\boldsymbol{1}^{M}||\boldsymbol{\nu})$
here, $\rho_{1}$ and $\rho_{2}$ represent the marginal regularization parameters, and $\mathrm{KL}(\boldsymbol{P}||\boldsymbol{Q})$ denotes the Kullback-Leibler divergence between two positive vectors. Similar to the classical OT formulation, there are solvers based on the Sinkhorn algorithm that can address Eq. (G.2)  (Pham et al., 2020). However, these solvers typically require more iteration steps to converge to optimal solutions due to the added complexity introduced by the relaxed marginal constraints.